OpenStudy (anonymous):

find the vertex, the line of symmetry, the maximum or minimum value of the quadratic function.. f(x)=-2x^2+2x+1

OpenStudy (anonymous):

the x value of the vertex is -b/2a = -2/-4 = 1/2. Plug that back into the equation to get the y. y=-2 * 1/4 + 2 * 1/4 + 1 = 1. So the vertex is at (1/2, 1). The parabola opens down the the vertex is a maximum , which equals 1. The line of symmetry is x=1/2.